# Comparing Apples to Oranges: A Dataset & Analysis of LLM Humour Understanding from Traditional Puns to Topical Jokes

### v1

@@ A.1 Evaluation Rubric
A.2 Implementation Details
@@
In order to select 150 jokes from the r/Jokes dataset to represent the topical subset. we sort the original dataset by the assigned ”score” field (calculated by Weller and Seppi (2020) based on the upvote-downvote ratio) to find the most popular jokes. Following this, we performed Named Entity Recognition (NER) via SpaCy, keeping jokes if they were found to contain entities of the following types from the SpaCy NER taxonomy: ORG (organisation), FAC (facility), WORK_OF_ART, LAW, ORG (organisations), PROD (products), EVENT, and PERSON. With these entity types consisting mainly of proper nouns of people, places and events, this filtering allows us to easily identify r/Jokes entries that are likely to require potentially esoteric world knowledge to perceive the humour within, as opposed to common sense or definitions of dictionary terms. This process reduced the dataset to just over 153k samples, of which we then take the top 150 highest scoring following manual verification.
@@
For all models, we used the following prompt: ”Explain the following joke (presented in square brackets) in approximately 100 words.nn[JOKE]nn”. This is based on good practice for clearly marking the boundaries between the task and the joke instance, whilst the 100-word length is based on the average length of the human-authored explanations.
@@
A.3 Dataset Enhancement
We additionally provide phonetic transcriptions in the International Phonetic Alphabet (IPA) of all jokes in the dataset. To do so, we utilise Anthropic’s Claude Sonnet 3.5 (20241022), owing to advanced handling of factors such as abbreviations over traditional grapheme-to-phoneme (G2P) models. A subset of these transcriptions were manually verified by an author of this work who is a trained phonetician. We additionally present a test phrase of ”The pikachu weighs 45 lbs and is 5ft 9”. He works at the CIA but previously worked at NASA”, in order to verify its handling of non-dictionary terms, abbreviations, initialisms and acronyms, which it did successfully. Specifically, we request transcriptions of the General American (GenAm) accent, owing to the primary userbase of Reddit with L1 English being North Americans.
@@ A.3 Dataset Enhancement
[…]
A.4 Additional Annotator & LLM Judge Results
@@
A.5 Additional Automatic Evaluation
@@
[…]
A.6 Additional Qualitative Analysis
@@
| 1 | ”What do you call a book club that’s been stuck on one book for years? Church.” |
| 2 | ”TIL unvaccinated children are less likely to be autistic. Because they are more likely to be dead.” |
| 3 | ”Set your wifi password to 2444666668888888. So when someone asks tell them it’s 12345678.” |
| 4 | ”My girlfriend is like the square root of -100. A solid 10, but also imaginary.” |
@@
| 1 | ”If America is storming Area 51 then the Europeans can storm the Vatican. We’ll take the aliens, you get the predators.” |
| 2 | ”For anyone attending Stan Lee’s funeral… Make sure you stay after the ceremony is finished.” |
| 3 | ”R Kelly is really changing the rap game. He takes the art out of rap artist.” |
| 4 | ”Obama smoked weed growing up, and now look where he is today. Unemployed with two kids and recently evicted.” |
Table 3 and Table 4 present examples of non-topical and topical jokes from the Reddit subsets, respectively. Regarding the non-topical jokes, the first example plays on common knowledge regarding religious practices and the focus of churches on the bible. Joke 2 is a form of deadpan humour that refers to the higher mortality rates of children who are not vaccinated against preventable diseases. The third joke presents phonetic wordplay, requiring the reader to interpret ”12345678” as ”one 2, 3 fours…” and so on. Finally, joke 4 plays on the mathematical concept of imaginary numbers. On the other hand, regarding the topical jokes, joke 1 makes reference to the media franchise of Alien v. Predator, whilst additionally referring to the 2019 news phenomena where a Facebook event was created to convince people to storm Area 51 to find aliens. The second joke refers to the Marvel film franchise and their penchant for extended scenes following the credits, in addition to Stan Lee’s real-life death. On the other hand, joke 3 makes light of a celebrity scandal involving R Kelly and sexual assault, performing wordplay that requires the letters ”art” to be removed from ”rap artist” to identify the particular scandal involving R Kelly. Finally, the fourth joke refers to Barack Obama losing his presidency to Donald Trump, and therefore being ”unemployed” and ”recently evicted” (from the White House).
@@
During the process of evaluating LLM responses, several unique response characteristics were observed. Firstly, the Google API refused to allow the Gemini models to produce explanations for some jokes, likely due to triggering content safeguards (5 explanations across 3 jokes, one of which the request was only blocked by Flash). Such instances were not notably any more/less offensive than other humour found within the Reddit joke subsets. In some instances, models provided textual responses regarding their refusal to address a particular joke (however, both cases only account for 1% of all jokes). Furthermore, in some instances, we observed that explanations contained subjective judgments as to the appropriateness of a particular joke, with the humour being deemed ”unacceptable”, without acknowledging the nuance and subjectivity of such a claim. Importantly, potentially offensive content was not exclusively found in the non-topical and topical Reddit joke subsets, with many of the traditional puns containing sexual innuendo and derogatory terms. Furthermore, whilst all of the models tested are capable of producing high-quality coherent text, we observed some instances of the Llama models resorting to degenerate patterns, such as being stuck in repetitive loops. Again, these account for a very small number of jokes (<1%). Regarding the heterographic pun subset of jokes, we noticed many instances of hallucination whereby the LLMs would explain that a punning word is phonetically similar to a word that shares little phonetic similarity. This is demonstrative of the challenge that heterographic puns present to LLMs, relying on phonetic characteristics of words, whilst such models are trained on orthographic text and therefore have limited phonetic knowledge.
A.7 LLM as a Judge Models and Prompts
@@
When selecting our model judge for automatic evaluation, we required that candidates not be from the same model family as any of those under evaluation or from the same family as each other. Our first model is the Mistral Mixture-of-Experts (Jiang et al., 2024) derivative, Prometheus 2 8x7B (Kim et al., 2024), which was specifically developed for LLM-based evaluation. The other two models are Qwen 72B Instruct (Qwen et al., 2025) and Gemma 2 9B Instruct (Riviere et al., 2024). These were selected as the highest performing models in the Hugging Face Judge Arena666https://huggingface.co/spaces/AtlaAI/judge-arena (with preliminary results disabled) that met our previously outlined criteria. At the time of selection (04/02/2025). Qwen was ranked 6th with 1203 ELO and Gemma 13th with 1157. The weights for the models were obtained from their respective Hugging Face model repositories. Overall, we found Qwen alone to demonstrate the higher alignment with our human evaluation (versus a jury ensemble or any other individual judge).
@@
Judge hyperparameters were all set to the model default except for temperature which was fixed to 0.1 for all models rather than zero based on prior work Sanh et al. (2022). To obtain scores from each judge, we used the following prompt. The scoring criteria was the same as given to the human annotators with the exception of score categories which were listed in ascending order rather than descending. This was done to better align with the training prompt of Prometheus.
@@
You are an expert judge evaluating explanations of jokes. You will be given a joke, a reference explanation (human-authored), and a model-generated explanation to evaluate.nn You must evaluate the model’s explanation on the matter of {criteria} using only the following scoring criteria:nn {scoring_criteria}nn The joke:n {joke}nn Reference explanation:n {reference_explanation}nn Model explanation to evaluate:n {model_explanation}nn Provide a score for this joke using only the scoring criteria provided. You must only respond with a single number between 0 and 5. You must produce no other output.

### v2

@@ 4.3 Evaluation Criteria
Accuracy entails whether or not the explanation contains correct material, and therefore does not negatively score content where parts of the joke have not been explained, but rather assesses the presence or absence of hallucinations and misunderstandings in a response. On the other hand, Completeness covers whether the fundamental elements of a joke have been addressed. Consequently, a joke may be fully explained (i.e., Completeness = 5), but additional information has been included which is incorrect (Accuracy 3). Conversely, an explanation may identify that a joke is a pun based on a specific word (i.e., Accuracy = 5), but omit explanations of the pun's meanings or other required information for understanding (i.e., Completeness 2).
Due to the significant amount of time required for rating 4800 explanations ( 40 hours), 1 author (who is a native speaker of English and holds degrees in Linguistics) rated all explanations. However, a subset of 320 explanations (10 jokes * 4 joke types * 8 models) was re-annotated by 2 third-party computer science students (who are also native English speakers).555Participants were given a 60 GBP Amazon voucher for approximately 3 hours of evaluation. These annotations were used exclusively to verify the primary evaluator's reliability. The results showed moderate levels of agreement for both accuracy ( = .574) and completeness ( = .553) using Krippendorff's alpha Krippendorff (2019), as well as high positive Pearson's correlations Pearson (1895) between the scores assigned by the primary human evaluator and third-party evaluators for accuracy (r = .765) and completeness (r = .795), both with . Overall, 82% of the Accuracy scores and 84% of the Completeness scores from the primary evaluator and the 2 third-party evaluators differed by 0 or 1 points, resulting in an 80.5% agreement between the main evaluator and third-party evaluators when converted to a binary pass/fail (as in 5.1). Across the whole dataset, we additionally leverage the LLM-as-a-judge paradigm using Qwen2.5-72B-Instruct, finding agreement for accuracy ( = .565) and completeness ( = .519) to be substantial, additionally supported by high Pearson's correlations for accuracy (r = .641) and completeness (r = .602), both with . See Appendix A.6 for details on the LLM-as-a-judge set up, and Appendix A.3 for a deeper analysis of the ratings from the third-party evaluators and LLM judge.
5 Human Evaluation
When comparing model performance, GPT-4o consistently outperforms all others, demonstrating the highest accuracy and completeness scores across joke types. Interestingly, GPT-4o Mini outperforms Gemini Pro in some specific cases, such as in accuracy on homographic jokes, as well as in both accuracy and completeness for non-topical and topical jokes. Additionally, open-source Llama 70B, outperforms most other models except GPT-4o, whilst the R1-based reasoning models underperform compared to most other non-reasoning models.
5.1 Explanation Success Rate
For an explanation to be useful, it must be both accurate and complete, rather than scoring highly on only one criterion. As a result, in Figure 4 we present a binary categorisation of explanations, treating scores of 4 or above on both criteria as being "good" quality explanations (i.e., containing only minimal, non-intrusive errors, and/or omitting minor details). This facilitates a clear comparison of the challenges different joke types pose.
As illustrated in Figure 4, the success rate of explanations varies significantly across different joke types. Homographic jokes, where a word has multiple meanings but identical spelling, consistently yield the highest proportion of successful explanations across nearly all models. This suggests that it is easier to navigate the ambiguity of homographs, likely due to language models being aware of the polysemy of the different pun words. In contrast, heterographic jokes, which involve words that sound the same but are spelt differently, pose more difficulty for all models. GPT-4o is the only model that maintains a relatively high success rate in this category. The challenge likely stems from the additional complexity introduced by needing to recognise the phonetic similarity of the pun word whilst being trained on orthographic tokens. Non-topical jokes also show varying levels of performance. The non-topical subset consists of jokes from Reddit, including puns, incongruities and situational irony that cover a wide range of joke forms that have been rated highly by the Reddit userbase, causing the LLMs to struggle more severely than traditional puns. Lastly, topical jokes, which require contextual awareness of specific events, present the most notable challenges. While GPT-4o generates a relatively high proportion of "good" explanations, other models are substantially less successful. This is likely due to the subtle references to pop culture and real-world entities in the "topical" subset presenting a significant challenge to current state-of-the-art LLMs, requiring both high levels of reasoning in addition to a form of information retrieval where the information that needs retrieving is not always explicitly referred to.
We see a unique pattern of performance from the Deepseek R1 models. Whilst R1 70B performs well overall, the model's reasoning capabilities do not result in it outperforming any other tested model. On the other hand, the 8B variant is the worst-performing model. We hypothesise that this is due to the focus on "reasoning" surrounding mathematical and coding problems, rather than "reasoning" in the general sense, such as identifying the common-sense incongruities that frequently cause humour, or "overthinking" the humour in simple puns. Furthermore, due to the reasoning steps taken, it may be the case that R1 8B begins with a moderately good answer (in line with the other models tested), but due to its uncertainty in explaining the jokes, the reasoning steps introduce many opportunities for hallucination, resulting in the smaller 8B model continuing to follow incorrect reasoning before presenting its final answer, therefore becoming progressively worse.
We perform a logistic regression analysis, which further supports our hypotheses. Model size has a strong and significant impact on explanation quality, with larger variants being far more likely to generate "good" explanations ( = 1.707, p < 0.001). This reinforces our finding that larger models significantly outperform smaller ones. Regarding joke difficulty, the results align with the predicted ordering. Homographic jokes are the easiest to explain ( = 0.583, p < 0.001), while non-topical ( = -0.511, p < 0.001) and topical jokes ( = -0.574, p < 0.001) are significantly harder. This confirms that jokes requiring external context pose a greater challenge to LLMs than those based on simple semantics and phonetics alone. Overall, the model is highly significant (p < 0.001) and explains a meaningful portion of the variance (pseudo R² = 0.138).
6 Automatic Evaluation
In addition to the previously presented human evaluation, we provide the results of automatic metrics in Table 1. Whilst existing reference-based automatic metrics are not well designed for this task, we do confirm our originally hypothesised ordering of joke difficulties, with the explanations of homographic/heterographic puns scoring higher than the non-topical joke subset, which in turn scores higher than the topical jokes subset on most metrics.
| Het. | Hom. | Non-Topical | Topical | |
| SacreBLEU | 8.53 | 10.15 | 7.86 | 7.09 |
| ROUGE-1 | 0.41 | 0.43 | 0.41 | 0.40 |
| ROUGE-2 | 0.12 | 0.15 | 0.13 | 0.12 |
| ROUGE-L | 0.25 | 0.28 | 0.25 | 0.23 |
| METEOR | 0.37 | 0.39 | 0.32 | 0.32 |
| BERTScore | 0.88 | 0.89 | 0.88 | 0.87 |
7 Case Study
Example LLM explanations of a randomly selected joke from the topical subset are shown in Figure 5. We omit the R1 reasoning models for space. The joke of interest refers to a 2018 internet phenomenon called the "Tide Pod Challenge", where teenagers would dare each other to consume toxic laundry pods (Tide being a specific brand). The central element to the joke's explanation is the suggestion that this new "teenage snack food market" increased sales, ergo generating enough revenue for the company to afford notoriously expensive advertising at the Super Bowl American football event. Interestingly, all of the full-size models inferred the reference to this challenge in their explanations. However, the smaller counterparts consistently omitted this information or presented misinterpretations in their explanations. GPT-4o Mini, for example, explains the joke by suggesting that "teenagers are known for being messy eaters", which caused increased sales. On the other hand, Llama 8B and Gemini Flash focus on the idea that the humour stems from "absurdity", without making any explicit mention of the real-world phenomenon, thereby failing to inform the user of important context. Further discussion of explanation characteristics can be found in Appendix A.5.
8 Discussion
Similarly, H2 concerned the hypothesis that homographic puns (which rely on polysemy) are easier to explain than heterographic puns, which was found to be the case in our tests. We observed that heterographic puns more frequently lead to hallucination (i.e., a lower accuracy score), owing to their reliance on phonetic characteristics, with LLMs having only implicit knowledge of how similar different words may "sound".
9 Conclusion
Limitations
In this work, we investigated the extent to which joke format and complexity affect the ability of LLMs to generate accurate and complete explanations. However, the nature of topical humour is that it is ever-evolving, with jokes relating to developing news stories and pop culture events continuously appearing. Consequently, our work focuses only on a subset of possible jokes and does not assess the ability of models to explain highly recent jokes, of which there may be only limited knowledge contained within a given model's training data (as the volume of reporting and discussion on a given topic grows over time).
Due to the jokes being from secondary sources (i.e., not novel jokes written for the purpose of this work), it is likely that the majority of the jokes have been seen before by the LLMs. However, explanations relating to these jokes are much less likely to be present. For the Reddit jokes in particular, the presence of explanations in the training data is reliant on individual Reddit users asking for an explanation within the relevant thread, and someone replying. Whilst this is possible, we sorted the r/Jokes dataset based on upvotes, and intuitively it is less likely that popular jokes require explanations, with Reddit's userbase already possessing the required knowledge in most cases (hence the jokes receiving a significant amount of upvotes). We manually verified a subset of the jokes and found readily available explanations to be rare.
Ethics Statement
We believe in and firmly adhere to the Code of Conduct in the performance of this work and the methods involved. Whilst the dataset element of our benchmark contains potentially offensive materials (reflecting the often controversial nature of humour found on Reddit), the aim of this resource is to offer explanations of potentially esoteric jokes in order to clarify where particular parties may be perceiving humour from. In this work, we do not create novel humour, nor present the included dataset as appropriate for training humour generation models. Whilst detoxifying and debiasing language models is a growing interest in the NLP community, humans are fundamentally opinionated and biased in ways that others may take offence at. Whilst we agree language models should not perpetuate these biases, the authors of this work believe that, if queried, language models should be able to make use of knowledge pertaining to human beliefs and opinions in offering impartial, objective explanations of phenomena such as humour. We therefore do not view the creation of an explanation for a joke to be equatable with an endorsement of said joke, but rather as an ideologically neutral task. We additionally consciously write our humour explanations in an objective manner, not lending unjust credence to unverified scandals or presenting opinions as facts. The study was improved by the internal ethics board of the lead author's institution, and all evaluators were provided with a particpant information sheet and consent form.
Acknowledgments
Tyler Loakman is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1].
Appendix A Appendix
A.1 Implementation Details
@@
In order to select 150 jokes from the r/Jokes dataset to represent the topical subset. we sort the original dataset by the assigned "score" field (calculated by Weller and Seppi (2020) based on the upvote-downvote ratio) to find the most popular jokes. Following this, we performed Named Entity Recognition (NER) via SpaCy, keeping jokes if they were found to contain entities of the following types from the SpaCy NER taxonomy: ORG (organisation), FAC (facility), WORK_OF_ART, LAW, ORG (organisations), PROD (products), EVENT, and PERSON. With these entity types consisting mainly of proper nouns of people, places and events, this filtering allows us to easily identify r/Jokes entries that are likely to require potentially esoteric world knowledge to perceive the humour within, as opposed to common sense or definitions of dictionary terms. This process reduced the dataset to just over 153k samples, of which we then take the top 150 highest scoring following manual verification.
@@
For all models, we used the following prompt: "Explain the following joke (presented in square brackets) in approximately 100 words.nn[JOKE]nn". This is based on good practice for clearly marking the boundaries between the task and the joke instance, whilst the 100-word length is based on the average length of the human-authored explanations.
@@
A.2 Dataset Enhancement
We additionally provide phonetic transcriptions in the International Phonetic Alphabet (IPA) of all jokes in the dataset. To do so, we utilise Anthropic's Claude Sonnet 3.5 (20241022), owing to advanced handling of factors such as abbreviations over traditional grapheme-to-phoneme (G2P) models. A subset of these transcriptions were manually verified by an author of this work who is a trained phonetician. We additionally present a test phrase of "The pikachu weighs 45 lbs and is 5ft 9". He works at the CIA but previously worked at NASA", in order to verify its handling of non-dictionary terms, abbreviations, initialisms and acronyms, which it did successfully. Specifically, we request transcriptions of the General American (GenAm) accent, owing to the primary userbase of Reddit with L1 English being North Americans.
@@ A.2 Dataset Enhancement
You are a linguistic transcription tool. For any text samples provided:nn 1. Convert EVERY sample into IPA format without exception - no samples should be skipped or partially transcribedn 2. Use General American (GenAm) pronunciation standards consistentlyn 3. If uncertain about any transcription, ask for clarification rather than making assumptionsn 4. Do not proceed without transcribing ALL text samples providednn Example input:n 'Hello world'nn Expected output:n h@”loU wldnn Please proceed with the transcription."
A.3 Additional Annotator & LLM Judge Results
@@
A.4 Additional Automatic Evaluation
@@
| METEOR | 0.25 | 0.28 | 0.24 | 0.23 | |
| BERTScore | 0.87 | 0.88 | 0.87 | 0.86 |
A.5 Additional Qualitative Analysis
@@
| 1 | "What do you call a book club that's been stuck on one book for years? Church." |
| 2 | "TIL unvaccinated children are less likely to be autistic. Because they are more likely to be dead." |
| 3 | "Set your wifi password to 2444666668888888. So when someone asks tell them it's 12345678." |
| 4 | "My girlfriend is like the square root of -100. A solid 10, but also imaginary." |
@@
| 1 | "If America is storming Area 51 then the Europeans can storm the Vatican. We’ll take the aliens, you get the predators." |
| 2 | "For anyone attending Stan Lee's funeral… Make sure you stay after the ceremony is finished." |
| 3 | "R Kelly is really changing the rap game. He takes the art out of rap artist." |
| 4 | "Obama smoked weed growing up, and now look where he is today. Unemployed with two kids and recently evicted." |
Table 3 and Table 4 present examples of non-topical and topical jokes from the Reddit subsets, respectively. Regarding the non-topical jokes, the first example plays on common knowledge regarding religious practices and the focus of churches on the bible. Joke 2 is a form of deadpan humour that refers to the higher mortality rates of children who are not vaccinated against preventable diseases. The third joke presents phonetic wordplay, requiring the reader to interpret "12345678" as "one 2, three 4…" and so on. Finally, joke 4 plays on the mathematical concept of imaginary numbers. On the other hand, regarding the topical jokes, joke 1 makes reference to the media franchise of Alien v. Predator, whilst additionally referring to the 2019 news phenomenon where a Facebook event was created to convince people to storm Area 51 to find aliens. The second joke refers to the Marvel film franchise and their penchant for extended scenes following the credits, in addition to Stan Lee's real-life death. On the other hand, joke 3 makes light of a celebrity scandal, involving wordplay that requires the sequence "art" to be removed from "rap artist" to identify a particular scandal involving R Kelly. Finally, the fourth joke refers to Barack Obama losing his presidency to Donald Trump, and therefore being "unemployed" and "recently evicted" (from the White House).
@@
During the process of evaluating LLM responses, several unique response characteristics were observed. Firstly, the Google API refused to allow the Gemini models to produce explanations for some jokes, likely due to triggering content safeguards (5 explanations across 3 jokes, where in one instance the request was only blocked by Flash). Such instances were not notably any more/less offensive than other humour found within the Reddit joke subsets. In some instances, models provided textual responses regarding their refusal to address a particular joke (however, both cases only account for 1% of all jokes). Furthermore, in some instances, we observed that explanations contained subjective judgments as to the appropriateness of a particular joke, with the humour being deemed "unacceptable", without acknowledging the nuance and subjectivity of such a claim. Importantly, potentially offensive content was not exclusively found in the non-topical and topical Reddit joke subsets, with many of the traditional puns containing sexual innuendo and derogatory terms. Furthermore, whilst all of the models tested are capable of producing high-quality coherent text, we observed some instances of the Llama models resorting to degenerate patterns, such as being stuck in repetitive loops. Again, these account for a very small number of jokes (<1%). Regarding the heterographic pun subset of jokes, we noticed many instances of hallucination whereby the LLMs would explain that a punning word is phonetically similar to a word that shares little phonetic similarity. This is demonstrative of the challenge that heterographic puns present to LLMs, relying on phonetic characteristics of words, whilst such models are trained on orthographic text and therefore have limited phonetic knowledge.
A.6 LLM-as-a-Judge Models and Prompts
@@
When selecting our LLM judge for automatic evaluation, we required that candidates not be from the same model family as any of those under evaluation or from the same family as each other. Our candidates included the Mistral Mixture-of-Experts (Jiang et al., 2024) derivative, Prometheus 2 8x7B (Kim et al., 2024), which was specifically developed for LLM-based evaluation, Qwen 72B Instruct (Qwen et al., 2025), and Gemma 2 9B Instruct (Riviere et al., 2024). These were selected due to being the best performing models in the Hugging Face Judge Arena666https://huggingface.co/spaces/AtlaAI/judge-arena (with preliminary results disabled) that met our previously outlined criteria. At the time of selection (04/02/2025), Qwen was ranked 6th with 1203 ELO and Gemma 13th with 1157. The weights for the models were obtained from their respective Hugging Face model repositories. Overall, we found Qwen alone to demonstrate the best alignment with our human evaluation (versus a jury ensemble or any other individual judge).
@@
Judge hyperparameters were all set to the model default except for temperature, which was fixed to 0.1 for all models, rather than zero, based on prior work Sanh et al. (2022). The scoring criteria were the same as given to the human annotators, with the exception of the ordering of score categories, which were listed in ascending order rather than descending. This was done to better align with the training prompt of Prometheus. To obtain scores from each judge, we used the following prompt:
@@
You are an expert judge evaluating explanations of jokes. You will be given a joke, a reference explanation (human-authored), and a model-generated explanation to evaluate.nn You must evaluate the model's explanation on the matter of {criteria} using only the following scoring criteria:nn {scoring_criteria}nn The joke:n {joke}nn Reference explanation:n {reference_explanation}nn Model explanation to evaluate:n {model_explanation}nn Provide a score for this joke using only the scoring criteria provided. You must only respond with a single number between 0 and 5. You must produce no other output.